solve the equation in exact differentials

2xydx+(x2-y2)dy=0



1
Expert's answer
2022-02-09T12:17:53-0500

Solution:

"2xydx+(x^2-y^2)dy=0
\\ \Rightarrow 2xydx=(y^2-x^2)dy
\\ \Rightarrow \dfrac{dy}{dx}=\dfrac{2xy}{y^2-x^2} \ ...(i)"

Put "y=vx"

"\dfrac{dy}{dx}=v+x\dfrac{dv}{dx}"

Using these in (i).

"v+x\dfrac{dv}{dx}=\dfrac{2vx^2}{v^2x^2-x^2}
\\ \Rightarrow x\dfrac{dv}{dx}=\dfrac{2v}{v^2-1}-v"

"\\ \Rightarrow x\dfrac{dv}{dx}=\dfrac{2v-v^3+v}{1-v^2}=\dfrac{3v-v^3}{1-v^2}
\\ \Rightarrow x\dfrac{dv}{dx}=\dfrac{v(3-v^2)}{1-v^2}
\\ \Rightarrow \dfrac{(1-v^2)dv}{v(3-v^2)}=\dfrac{dx}{x}"

"\\ \Rightarrow \dfrac{1}{v\left(3-v^2\right)}dv- \dfrac{v^2}{v\left(3-v^2\right)}dv=\dfrac{dx}{x}"

Using partial fractions,

"[\dfrac{1}{3v}-\dfrac{1}{6\left(v+\sqrt{3}\right)}-\dfrac{1}{6\left(v-\sqrt{3}\right)}]dv-\dfrac{v}{3-v^2}dv=\dfrac{dx}{x}
\\ \Rightarrow [\dfrac{1}{3v}-\dfrac{1}{6\left(v+\sqrt{3}\right)}-\dfrac{1}{6\left(v-\sqrt{3}\right)}]dv+\dfrac12.\dfrac{-2v}{3-v^2}dv=\dfrac{dx}{x}"

On integrating,

"\frac{1}{3}\ln \left|v\right|-\frac{1}{6}\ln \left|v+\sqrt{3}\right|-\frac{1}{6}\ln \left|v-\sqrt{3}\right|-\left(-\frac{1}{2}\ln \left|3-v^2\right|\right)=\ln |x|+\ln C"

"\Rightarrow \frac{1}{3}\ln \left|\frac yx\right|-\frac{1}{6}\ln \left|\frac yx+\sqrt{3}\right|-\frac{1}{6}\ln \left|\frac yx-\sqrt{3}\right|+\frac{1}{2}\ln \left|3-\frac {y^2}{x^2}\right|=\ln |Cx|
\\ \Rightarrow \ln |\frac yx|^{1/3}-\ln |\frac yx+\sqrt3|^{1/6}-\ln|\frac yx-\sqrt3|^{1/6}+\ln|3-\frac {y^2}{x^2}|^{1/2}=\ln|Cx|
\\ \Rightarrow \ln |\dfrac{(\frac yx)^{1/3}(3-\frac {y^2}{x^2})^{1/2}}{(\frac yx+\sqrt3)^{1/6}(\frac yx-\sqrt3)^{1/6}}|=\ln |Cx|"

"\Rightarrow \ln |\dfrac{(\frac yx)^{1/3}(3-\frac {y^2}{x^2})^{1/2}}{(3-\frac {y^2}{x^2})^{1/6}}|=\ln |Cx|
\\ \Rightarrow \ln |{(\frac yx)^{1/3}(3-\frac {y^2}{x^2})^{1/3}}|=\ln |Cx|
\\ \Rightarrow {(\frac yx)^{1/3}(3-\frac {y^2}{x^2})^{1/3}}= Cx
\\ \Rightarrow (\frac yx)(3-\frac {y^2}{x^2})=Kx^3 \ \ [\because K=C^3]
\\ \Rightarrow 3yx^2-y^3=Kx^6"